1/(w+a)+1/(w+b)+1/(w+c)=2w2


and


1/(w2+a)+1/(w2+b)+1/(w2+c)=2w


what is


1/(a+1)+1/(b+1)+1/(c+1)=


A)2


B)-2


C)w2


D)w


where w is a complex cube root of unity.


 

Grade:

4 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear srichaitanya

let f(z) =1/(z+a)+1/(z+b)+1/(z+c)  a function in z

    given f(w) =2w2

and        f(w2)=2w

clearly one of the possible function is

    f(z)=2z2

 

so f(1)=2


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best
srichaitanya.
 
Regards,
Askiitians Experts
Badiuddin

mycroft holmes
14 years ago

I wonder what is the basis for assuming that f(z) = z2 here.

 

Multiplying the first relation by w, we get

 

w/w+a + w/w+b + w/w+c = 2

 

or a/w+a + b/w+b + c/w+c = 1

 

Similarly from the second relation, we get a/a+w2 + b/b+w2 + c/c+w2 = 1

 

Thus, w and w2 are the roots of a/(x+a) + b/(x+b) + c/(x+c) = 1

 

which simplifies to x3 - (ab+bc+ca)x -2abc = 0. This means that the sum of roots of this cubic is zero. Let the third root be r.

 

Then r + w+w2 = 0 which means r = 1.

 

Thus a/a+1 + b/b+1 + c/c+1 = 1 or 1/a+1 + 1/b+1 + 1/c+1 = 2

Ass satwik
5 years ago
Can you expalex me the step 
a/a+w + b/b+w + c/c+w = 1
Please.........
.............
...........
..................
.........
 
 
 
 
 
 
 
 
Allison
5 years ago
W3 = 1 therefore w x w​= 1 therefore w​= 1/w --(1)
1/a+w + 1/b+w +  1/c+w = 2w​= 2/w. (From 1)
 
Now take conjugate of the above entire equation, we get (I am representing conjugate of w as “p" for convenience)
1/a+p + 1/b+p + 1/c+p = 2/p
Now w is -1/2 +root3/2 and w²  is -1/2 – root 3/2 which is just the same as conjugate of w
Therefore w² =p
So,
1/ a+w2 + 1/b+w2 + 1/c+w2 = 2/w2
So w and w2 are riits of equation 1/a+x + 1/b+x + 1/c+x = 2/x
Nowhow puthow x=1 to get the answer as 2.